Valdosta Chemical Company manufactures two industrial chemical products in a joint process. In May, 16,000 gallons of input costing $63,000 were processed at a cost of $165,000. The joint process resulted in 12,000 pounds of Resoline and 4,000 pounds of Krypto. Resoline sells for $25 per pound, and Krypto sells for $50 per pound. Management generally processes each of these chemicals further in separable processes to produce more refined chemical products. Resoline is processed separately at a cost of $6 per pound. The resulting product, Resolite, sells for $33 per pound. Krypto is processed separately at a cost of $15 per pound. The resulting product, Kryptite, sells for $69 per pound.

Required:

  1. 2-a. Allocate the company’s joint production costs for May using the physical-units method.

  2. 2-b. Allocate the company’s joint production costs for May using the relative-sales-value method.

  3. 2-c. Allocate the company’s joint production costs for May using the net-realizable-value method.

  4. 3-a. Valdosta’s management is considering an opportunity to process Kryptite further into a new product called Omega. The separable processing will cost $38 per pound. Packaging costs for Omega are projected to be $7 per pound, and the anticipated sales price is $105 per pound. Calculate the incremental profit or loss from processing Kryptite into Omega.

  5. 3-b. Should Kryptite be processed further into Omega?
Answer #1

2-a. Allocation of joint costs using physical units method

Resoline = (63,000+165,000)*12,000/16,000 = $171,000

Krypto = (63,000+165,000)*4,000/16,000 = $57,000

2-b Using relative sales value

Resoline = 228,000*(12,000*25)/(12,000*25+4,000*50)

= $136,800

Krypto = 228,000*200,000/500,000 = $91,200

2-c Using NRV Method

Resoline = 228,000*(12,000*27)/(12,000*27+4,000*54)

= $136,800

Krypto = $91,200

3-a. Incremental Profit = (105-38-7-69)*4,000

= -$36,000

i.e. Loss of $36,000

Note: Costs incurred upto Kryptite are not relevant for this decision since it will be incurred whether or not it is further processed into Omega

Kryptite are not relevant for this decision since it will be incurred whether or not it is further processed into Omega

3-b. No, Kryptite Should not be processed further into Omega since it will reduce income